Please show formulas. Jackson Company is considering a project that has the following cash flows. What are the project's payback, discounted payback, NPV, IRR, and MIRR? The weighted Average Cost of capital of Jackson Company is 10 percent. Explain, in writing, if the project is accepted and why?

0

-950

1

525

2

485

3

445

4

405
Answer #1

Please refer to below spreadsheet for calculation and answer. Cell reference also provided.

C D E F G 1 2 Cash Flow ($950.00) $525.00 $485.00 $445.00 $405.00 10% Cumulative CF ($950.00) ($425.00) $60.00 $505.00 $910.0

Cell reference -

B C D E F G Year Cash Flow -950 525 485 445 405 0.1 Cumulative CF PVF(10%) PV CF Cumulative PV CF =C3 =1/1.1^B3 =C3*E3 =F3 =D

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.